# Scaling LLM Planning: NL2FLOW for Parametric Problem Generation and Rigorous Evaluation

### v1

@@ 5 Conclusion
This study introduced a fully automated pipeline for generating and evaluating reasoning data, leveraging a planning data generator capable of producing both PDDL and intermediate Python object representations alongside a symbolic evaluator for assessing plan validity and JSON translation accuracy. This approach enables scalable data creation and rigorous assessment crucial for advancing large language models with reasoning and planning capabilities.
My evaluation, using simple planning problems generated by NL2FLOW, revealed that current open-sourced, instruct-tuned LLMs demonstrate a surprising capacity for direct plan generation. Notably, I observed that the highest success rate for translating natural language into a JSON representation of a plan was lower than the highest rate of generating a valid plan directly. This suggests that unnecessarily decomposing the reasoning task – breaking it down into intermediate translation steps – may actually hinder performance, implying a benefit to models learning to reason directly from natural language to action. Ultimately, the bottlenecks and sources of error within a reasoning system are likely to shift depending on the complexity and scale of the problem. As I push the boundaries of LLM reasoning, understanding these dynamic limitations will be paramount.
Further analysis through regression revealed that the impact of problem characteristics on plan generation is significantly dependent on the specific language model and the prompting strategy employed. This highlights the critical need for model-specific insights; what works for one LLM may not generalize to another.

### v2

@@ 5 Conclusion
A key contribution of this work is the introduction of a fully automated pipeline for generating and evaluating reasoning data. This pipeline, capable of producing intermediate Python object representations, natural language descriptions, and PDDLs, coupled with a symbolic workflow evaluator and JSON translation accuracy assessment, offers a robust methodology for systematically probing the limitations and boundaries of LLMs as workflow generators (planners).
This study showed that translating natural language into an intermediate JSON workflow problem representation did not improve overall workflow generation performance best (Figure 1); in fact, it yielded lower success rates than direct plan generation. This suggests that decomposing the reasoning task into multiple translation steps can introduce unnecessary complexity, highlighting the potential benefit of direct reasoning from language to action. Meanwhile, as LLM reasoning scales to increasingly complex problems, the bottlenecks and sources of error within AI planning systems can shift.
This work demonstrates that current open-source, instruct-tuned LLMs can generate effective plans directly from natural language instructions, outperforming previous LLM planning studies that relied on formal PDDL representations [20, 21, 22]. This difference likely arises from the nature of the reasoning tasks: our study focused on workflow creation, where generated actions inherently involve fixed input and output variables, contrasting with the flexible variable binding common in domains like the Blocks World. Importantly, the complexity of workflow creation is comparable to that of domains used in prior LLM planning research. The observed pattern reinforces the hypothesis that LLM planning capabilities are domain-specific, potentially rooted in their statistical reliance on token prediction and the associated sensitivity to the compositional complexity of input tokens.
Further analysis through regression confirmed that the impact of problem characteristics on plan generation is contingent upon both the specific language model and the prompting strategy used. This underscores the critical need for model-specific and prompt-specific insights; generalizations across LLMs should be approached with caution.
NL2Flow extends beyond simple LLM assessment by providing a framework for evaluating the planning capabilities of novel agentic architectures. These architectures vary in agent and tool configurations and operate with randomly assigned goals, expressed as LLM-generated utterances based on system-defined agent and tool descriptions (including those adhering to the Model Context Protocol (MCP)). By translating these goals and configurations into PDDL, NL2Flow enables the evaluation of agentic system trajectories using a standardized planning framework, allowing for assessment of an agent’s ability to select and compose tools based on custom, symbolically-verifiable workflows. Given the prevalence of Chain-of-Thought (CoT) prompting, to decompose tasks into manageable steps for workflow creation, in agentic systems for workflow creation, verifying the soundness of this semantic approach is crucial for improving system reliability.
Future work will focus on expanding this automated pipeline to generate datasets exhibiting greater combinatorial complexity and incorporating diverse constraints. Investigating the interplay between prompting strategies, model architecture, and dataset characteristics will be crucial for understanding how to best use LLMs for complex workflow automation. Ultimately, a systematic analysis of LLM limitations across a spectrum of reasoning tasks will be vital for guiding future development and realizing the full potential of these models.